# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ III.3 Proton–proton collisions
Another important measurement related to the enhancement of the $\Lambda_{\mathrm{c}}^{+}$ baryon performed by the ALICE Collaboration, is the measurement of the $\Sigma_{\mathrm{c}}^{0,+,++}$ baryon at midrapidity in pp collisions at $\sqrt{s}$ = 13 $\mathrm{TeV}$ [176]. As discussed in Sect. III.1, the production of $\Sigma_{\mathrm{c}}^{0,+,++}$ states is suppressed with respect to $\Lambda_{\mathrm{c}}^{+}$-states [103] in $\mathrm{e^{+}e^{-}}$ collisions. As shown in the top-right panel of Fig. 13, in pp collisions the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio is close to 0.2 for 2 $<$ $p_{\mathrm{T}}$ $<$ 6 GeV$/c$, and decreases with $p_{\mathrm{T}}$ down to about 0.1 for 8 $<$ $p_{\mathrm{T}}$ $<$ 12 GeV$/c$, though the uncertainties do not allow firm conclusions about the $p_{\mathrm{T}}$ dependence. From Belle measurements (Table 4 in Ref. [103]), the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\Lambda_{\mathrm{c}}^{+}$ ratio in $\mathrm{e^{+}e^{-}}$ collisions at $\sqrt{s}$ = 10.52 GeV can be evaluated to be around 0.17 and, thus, the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio can be estimated to be around 0.02. Therefore, a remarkable difference is present between pp and $\mathrm{e^{+}e^{-}}$ collision systems. In Ref. [176] the $\Sigma_{\mathrm{c}}^{0,+,++}$ feed-down fraction for $\Lambda_{\mathrm{c}}^{+}$ with $2<p_{\mathrm{T}}<$ $\mathrm{GeV}/c$ was measured to be $\Lambda_{\mathrm{c}}^{+}\leftarrow\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}=\mathrm{0.38\pm 0.06~{}(stat)\pm 0.06~{}(syst)}$, a significantly larger value than the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\Lambda_{\mathrm{c}}^{+}$ ratio measured by Belle and the $\sim$0.13 expectation from PYTHIA 8 Monash tune simulations. This indicates a larger increase of the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio than of the direct-$\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio from $\mathrm{e^{+}e^{-}}$ to pp collisions. The larger feed-down from $\Sigma_{\mathrm{c}}^{0,+,++}$ partially explains the difference between the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios in pp and $\mathrm{e^{+}e^{-}}$ collisions. It suggests that the suppression of $\Sigma_{\mathrm{c}}^{0,+,++}$ states observed in $\mathrm{e^{+}e^{-}}$ collisions, which in the Lund string-fragmentation picture is ascribed to the lower production yield of spin-1 than spin-0 diquarks in string breaking, is absent or reduced in pp collisions. In the models that provide a fair description of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio in pp collisions this suppression mechanism is absent or heavily reduced, and a sizeable contribution to $\Lambda_{\mathrm{c}}^{+}$ production from strong decays of $\Sigma_{\mathrm{c}}^{0,+,++}$ states is expected. As shown in Fig. 13, the QCD-based CR-BLC (for which the three variations defined in Ref. [33] are considered), SHM+RQM, and Catania models describe, within uncertainties, both the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. In the SHM, only the hadron mass and spin degeneracy factor determine the hadron yields. In all these models the production of $\Sigma_{\mathrm{c}}^{0,+,++}$ (2455) is expected to increase by significant factors in pp with respect to $\mathrm{e^{+}e^{-}}$ collisions and to become comparable or even larger than that of direct $\Lambda_{\mathrm{c}}^{+}$. In Fig. 14 it is illustrated how the current measurements can provide useful constraints for a better tuning of the PYTHIA 8 QCD-based CR-BLC parameters relating to the production of spin-1 compared to spin-0 charm-light diquarks. When the charm-light diquark is coupled to a quark to form a baryon, the coupling in $l=0$ state of a diquark with spin=0 cannot result in baryons with spin=3/2, which are instead possible with spin=1 diquarks. Excited $\Lambda_{\mathrm{c}}^{+}$ and $\Sigma_{\mathrm{c}}$ states with spin=3/2 decay preferentially to $\Lambda_{\mathrm{c}}^{+}$ rather than to $\Sigma_{\mathrm{c}}^{0,+,++}$(2455) [49]. Therefore, increasing the probability of producing spin=1 charm-light diquarks changes the feed-down contribution to $\Sigma_{\mathrm{c}}^{0,+,++}$ and $\Lambda_{\mathrm{c}}^{+}$, effectively resulting in a reduction of the $\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}$ ratio. Thus one can significantly improve the description of the measured feed-down contribution from $\Sigma_{\mathrm{c}}^{0,+,++}$ to $\Lambda_{\mathrm{c}}^{+}$ production in PYTHIA by increasing the probability of spin-1 charm-light diquark formation in junctions. In PYTHIA 8, this probability is regulated by the parameter probQQ1toQQ0join, which dictates the suppression of forming a junction diquark with spin-1 compared to spin-0, acting on top of spin-state counting that dictates a 3-to-1 enhancement. The parameter probQQ1toQQ0join allows for different suppression factors given the heaviest flavour involved in the junction diquark formation is up/down, strange, charm or bottom. A value close to zero implies a large suppression of spin-1 junction diquarks, while probQQ1toQQ0join=1 implies no suppression. Though the data uncertainties do not allow a fine tuning of the charm probQQ1toQQ0join parameter, values around 0.5 or higher seem preferable as seen in Fig. 14. Note this is in contrast to the parameter probQQ1toQQ0 which determines the spin-1 to spin-0 probability for diquark-antidiquark production from standard string breaks, which has a default LEP-tuned value of 0.0275. This however is not entirely surprising as diquark-antidiquark pair creation must be light-flavoured, and thus probQQ1toQQ0 will only effect light diquark production. Such a large suppression of spin-1 light diquarks can be justified by mass difference arguments and the smaller binding energy of spin-1 diquarks. Contrastingly in a heavy-light diquark that is formed in a junction, one would expect the spin splitting to almost decouple. Nonetheless there remains large ambiguity in probQQ1toQQ0join in the light-sector, and probQQ1toQQ0join values for all flavours were not well constrained for the default PYTHIA values nor the QCD-based CR tunes. Future measurements of the production of excited states, in particular of $\Lambda_{\mathrm{c}}^{+}(2595)$, $\Lambda_{\mathrm{c}}^{+}(2625)$ and $\Sigma_{\mathrm{c}}^{0.+.++}(2520)$ will be useful to constrain this parameter, as illustrated in Fig. 14 below.
@@
The ALICE Collaboration extended the measurement of charm-baryon production to baryons with strange quarks among the constituent quarks, measuring the $\Xi_{\mathrm{c}}^{+,0}$ cross section in pp collisions at $\sqrt{s}$ = 5.02 TeV [179] and $\sqrt{s}$ = 13 TeV [180] and the $\Omega_{\mathrm{c}}^{0}$ [177] cross section times branching ratio in pp collisions at $\sqrt{s}$ = 13 TeV. The bottom left panel of Fig. 13 shows the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratios measured as a function of $p_{\mathrm{T}}$ at midrapidity in pp collisions at $\sqrt{s}$ = 13 TeV. The values observed are generally lower than those of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio, but the $p_{\mathrm{T}}$ dependence is similar. The PYTHIA 8 Monash tune significantly underestimates the data by a factor of 23–43 in the low-$p_{\mathrm{T}}$ region and by a factor of about 5 in the highest $p_{\mathrm{T}}$ interval. All three CR modes give a similar magnitude and $p_{\mathrm{T}}$-dependence of $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and predict a larger ratio with respect to the Monash tune. However, differently from what is observed for the $\Lambda_{\mathrm{c}}^{+}$ and the $\Sigma_{\mathrm{c}}^{0,+,++}$ baryons, they underestimate the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio by a factor 4–9 for $p_{\mathrm{T}}$ $<$4 GeV$/c$. At higher $p_{\mathrm{T}}$ the measured $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio is closer to PYTHIA 8 calculations. As the CR-BLC models underpredict the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio despite well describing $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$, the source of this discrepancy is primarily assumed to be the strange content of the $\Xi_{\mathrm{c}}^{+,0}$ baryon, where we suspect strangeness enhancement effects particularly at low $p_{\mathrm{T}}$. There have been some models such as Rope hadronization [47] which aim to implement a strangeness enhancement mechanism, however the rope model thus far has not been able to describe both the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ ratios simultaneously (or equivalently stated, it cannot describe the $\Xi_{\mathrm{c}}^{+,0}$/$\Lambda_{\mathrm{c}}^{+}$ ratio). The SHM+RQM model [38], which captures the magnitude and $p_{\mathrm{T}}$ dependence of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ yield ratios, also underestimates the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio. The measured ratios are also compared to the QCM [35] and the Catania [36, 37] models that include hadronization via coalescence and reproduce well the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. They both tend to underestimate the data though, considering the uncertainties, the Catania model remains close to the measured values over the whole $p_{\mathrm{T}}$ interval. Another interesting observation is given by the fact that the $\Xi_{\mathrm{c}}^{0}$/$\Sigma_{\mathrm{c}}^{0,+,++}$ ratio shown in Ref [180] is described well by the default PYTHIA 8 Monash tune [11], which significantly underestimates both $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. Though this could be accidental (the predicted value remains to be explained along with its dependence on the two almost identical particle masses, and the assumed feed-down contributions), it may also signal the removal of a similar suppression factor affecting both $\Xi_{\mathrm{c}}^{0}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$ production in $\mathrm{e^{+}e^{-}}$ collisions. In particular, in most calculations [134], a similar mass is obtained or assumed for S=1 (ud,uu,dd) diquarks as contained in $\Sigma_{\mathrm{c}}^{0,+,++}$ and S=0 (ds) diquarks as in $\Xi_{\mathrm{c}}^{0}$. As discussed for the $\Lambda_{\mathrm{c}}^{+}\leftarrow\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}$ ratio, this could suggest that light diquarks are more easily produced in pp than $\mathrm{e^{+}e^{-}}$ collisions via other processes than the Schwinger mechanism (e.g. colour reconnections or coalescence), or that in pp collisions charm-baryon formation does not proceed via the combination of a charm quark with a light diquark. As mentioned in Sect. [sec:theoryCoalescence], the coalescence mechanism can be factorized in terms of a formation of diquark recombining with a further quark. In the limit of negligible binding energy this would correspond to the coalescence model already discussed. However, this is different from the assumption of the existence of a thermal yield of diquarks with masses that are not simply the sum of the two quark masses. In this respect it is also argued in Ref. [181] that a further enhancement of $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio could be due to the large attraction of the $us$ and $ds$ diquark channel with respect to the $ud$ one which is suggested by a relativistic quark model with a potential having a large screened mass as extracted from recent LQCD data. Introducing such an additional coalescence component in the Catania model could strongly improve its agreement with experimental data  [181], but within the current experimental data uncertainties no conclusion can be drawn. Clarifying the role of light and heavy diquarks will be important for understanding hadronization in pp as well as in heavy-ion collisions, where diquarks could also be effective degrees of freedom of the system close to the phase transition [96].

### v3

@@ III.3 Proton–proton collisions
Another important measurement related to the enhancement of the $\Lambda_{\mathrm{c}}^{+}$ baryon performed by the ALICE Collaboration, is the measurement of the $\Sigma_{\mathrm{c}}^{0,+,++}$ baryon at midrapidity in pp collisions at $\sqrt{s}$ = 13 $\mathrm{TeV}$ [18]. As discussed in Sect. III.1, the production of $\Sigma_{\mathrm{c}}^{0,+,++}$ states is suppressed with respect to $\Lambda_{\mathrm{c}}^{+}$-states [120] in $\mathrm{e^{+}e^{-}}$ collisions. As shown in the top-right panel of Fig. 13, in pp collisions the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio is close to 0.2 for 2 $<$ $p_{\mathrm{T}}$ $<$ 6 GeV$/c$, and decreases with $p_{\mathrm{T}}$ down to about 0.1 for 8 $<$ $p_{\mathrm{T}}$ $<$ 12 GeV$/c$, though the uncertainties do not allow firm conclusions about the $p_{\mathrm{T}}$ dependence. From Belle measurements (Table 4 in Ref. [120]), the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\Lambda_{\mathrm{c}}^{+}$ ratio in $\mathrm{e^{+}e^{-}}$ collisions at $\sqrt{s}$ = 10.52 GeV can be evaluated to be around 0.17 and, thus, the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio can be estimated to be around 0.02. Therefore, a remarkable difference is present between pp and $\mathrm{e^{+}e^{-}}$ collision systems. In Ref. [18] the $\Sigma_{\mathrm{c}}^{0,+,++}$ feed-down fraction for $\Lambda_{\mathrm{c}}^{+}$ with $2<p_{\mathrm{T}}<$ $\mathrm{GeV}/c$ was measured to be $\Lambda_{\mathrm{c}}^{+}\leftarrow\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}=\mathrm{0.38\pm 0.06~{}(stat)\pm 0.06~{}(syst)}$, a significantly larger value than the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\Lambda_{\mathrm{c}}^{+}$ ratio measured by Belle and the $\sim$0.13 expectation from PYTHIA 8 Monash tune simulations. This indicates a larger increase of the $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratio than of the direct-$\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio from $\mathrm{e^{+}e^{-}}$ to pp collisions. The larger feed-down from $\Sigma_{\mathrm{c}}^{0,+,++}$ partially explains the difference between the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratios in pp and $\mathrm{e^{+}e^{-}}$ collisions. It suggests that the suppression of $\Sigma_{\mathrm{c}}^{0,+,++}$ states observed in $\mathrm{e^{+}e^{-}}$ collisions, which in the Lund string-fragmentation picture is ascribed to the lower production yield of spin-1 than spin-0 diquarks in string breaking, is absent or reduced in pp collisions. In the models that provide a fair description of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio in pp collisions this suppression mechanism is absent or heavily reduced, and a sizeable contribution to $\Lambda_{\mathrm{c}}^{+}$ production from strong decays of $\Sigma_{\mathrm{c}}^{0,+,++}$ states is expected. As shown in Fig. 13, the QCD-based CR-BLC (for which the three variations defined in Ref. [51] are considered), SHM+RQM, and Catania models describe, within uncertainties, both the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. In the SHM, only the hadron mass and spin degeneracy factor determine the hadron yields. In all these models the production of $\Sigma_{\mathrm{c}}^{0,+,++}$ (2455) is expected to increase by significant factors in pp with respect to $\mathrm{e^{+}e^{-}}$ collisions and to become comparable or even larger than that of direct $\Lambda_{\mathrm{c}}^{+}$. In Fig. 14 it is illustrated how the current measurements can provide useful constraints for a better tuning of the PYTHIA 8 QCD-based CR-BLC parameters. Specifically, in junction systems with slow heavy quarks, charm-light diquarks are introduced in an intermediate step during the fragmentation process (see [192] for details). These should probably not be thought of as a physical representation of the internal structure of the baryon being formed (see the discussion in section III.1), but mainly as a technical bookkeeping device. Nevertheless these diquarks formally need to be assigned a spin quantum number, 0 or 1. When the charm-light diquark is combined with a quark to form a baryon, the spin assigned to the charm-light diquark is imprinted on the rate of spin-3/2 baryons since spin=3/2 can only be produced (in an $l=0$ state) if the charm-light diquark has spin=1, not otherwise.
In PYTHIA 8, this probability is regulated by the parameter probQQ1toQQ0join, which dictates the suppression of forming a junction diquark with spin-1 compared to spin-0, acting on top of spin-state counting that dictates a 3-to-1 enhancement. The parameter probQQ1toQQ0join allows for different suppression factors given the heaviest flavour involved in the junction diquark formation is up/down, strange, charm or bottom. A value close to zero implies a large suppression of spin-1 junction diquarks, while probQQ1toQQ0join=1 implies no suppression. Given the argument above that the charm-light diquark should probably not be attributed significant physical meaning, it would be reasonable to expect this parameter not to differ greatly from unity.
Excited $\Lambda_{\mathrm{c}}^{+}$ and $\Sigma_{\mathrm{c}}$ states with spin=3/2 decay preferentially to $\Lambda_{\mathrm{c}}^{+}$ rather than to $\Sigma_{\mathrm{c}}^{0,+,++}$(2455) [68]. Therefore, increasing the probability of producing spin=1 charm-light diquarks changes the feed-down contribution to $\Sigma_{\mathrm{c}}^{0,+,++}$ and $\Lambda_{\mathrm{c}}^{+}$, effectively resulting in a reduction of the $\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}$ ratio. Thus one can significantly improve the description of the measured feed-down contribution from $\Sigma_{\mathrm{c}}^{0,+,++}$ to $\Lambda_{\mathrm{c}}^{+}$ production in PYTHIA by increasing the probability of spin-1 charm-light diquark formation in junctions.
Though the data uncertainties do not allow a fine tuning of the charm probQQ1toQQ0join parameter, values around 0.5 or higher seem preferable as seen in Fig. 14. Note this is in contrast to the parameter probQQ1toQQ0 which determines the spin-1 to spin-0 probability for light diquark-antidiquark production from standard string breaks, which has a default LEP-tuned value of 0.0275. Such a large suppression of spin-1 light diquarks can be justified by mass difference arguments and the smaller binding energy of spin-1 diquarks. Nonetheless there remains large ambiguity in probQQ1toQQ0join in the light sector, and probQQ1toQQ0join values for all flavours were not well constrained for the default PYTHIA values nor the QCD-based CR tunes. Future measurements of the production of excited states, in particular of $\Lambda_{\mathrm{c}}^{+}(2595)$, $\Lambda_{\mathrm{c}}^{+}(2625)$ and $\Sigma_{\mathrm{c}}^{0.+.++}(2520)$ will be useful to constrain this parameter, as illustrated in Fig. 14 below.
@@
The ALICE Collaboration extended the measurement of charm-baryon production to baryons with strange quarks among the constituent quarks, measuring the $\Xi_{\mathrm{c}}^{+,0}$ cross section in pp collisions at $\sqrt{s}$ = 5.02 TeV [20] and $\sqrt{s}$ = 13 TeV [21] and the $\Omega_{\mathrm{c}}^{0}$ [22] cross section times branching ratio in pp collisions at $\sqrt{s}$ = 13 TeV. The bottom left panel of Fig. 13 shows the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratios measured as a function of $p_{\mathrm{T}}$ at midrapidity in pp collisions at $\sqrt{s}$ = 13 TeV. The values observed are generally lower than those of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ ratio, but the $p_{\mathrm{T}}$ dependence is similar. The PYTHIA 8 Monash tune significantly underestimates the data by a factor of 23–43 in the low-$p_{\mathrm{T}}$ region and by a factor of about 5 in the highest $p_{\mathrm{T}}$ interval. All three CR modes give a similar magnitude and $p_{\mathrm{T}}$-dependence of $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and predict a larger ratio with respect to the Monash tune. However, differently from what is observed for the $\Lambda_{\mathrm{c}}^{+}$ and the $\Sigma_{\mathrm{c}}^{0,+,++}$ baryons, they underestimate the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio by a factor 4–9 for $p_{\mathrm{T}}$ $<$4 GeV$/c$. At higher $p_{\mathrm{T}}$ the measured $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio is closer to PYTHIA 8 calculations. As the CR-BLC models underpredict the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio despite well describing $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$, the source of this discrepancy is primarily assumed to be the strange content of the $\Xi_{\mathrm{c}}^{+,0}$ baryon, where we suspect strangeness enhancement effects particularly at low $p_{\mathrm{T}}$. There have been some models such as Rope hadronization [65] which aim to implement a strangeness enhancement mechanism, however the rope model thus far has not been able to describe both the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and $\Lambda_{\mathrm{c}}^{+}$/$\mathrm{D^{0}}$ ratios simultaneously (or equivalently stated, it cannot describe the $\Xi_{\mathrm{c}}^{+,0}$/$\Lambda_{\mathrm{c}}^{+}$ ratio). The SHM+RQM model [56], which captures the magnitude and $p_{\mathrm{T}}$ dependence of the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ yield ratios, also underestimates the $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio. The measured ratios are also compared to the QCM [53] and the Catania [54, 55] models that include hadronization via coalescence and reproduce well the $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. They both tend to underestimate the data though, considering the uncertainties, the Catania model remains close to the measured values over the whole $p_{\mathrm{T}}$ interval. Another interesting observation is given by the fact that the $\Xi_{\mathrm{c}}^{0}$/$\Sigma_{\mathrm{c}}^{0,+,++}$ ratio shown in Ref [21] is described well by the default PYTHIA 8 Monash tune [11], which significantly underestimates both $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$/$\mathrm{D^{0}}$ ratios. Though this could be accidental (the predicted value remains to be explained along with its dependence on the two almost identical particle masses, and the assumed feed-down contributions), it may also signal the removal of a similar suppression factor affecting both $\Xi_{\mathrm{c}}^{0}$ and $\Sigma_{\mathrm{c}}^{0,+,++}$ production in $\mathrm{e^{+}e^{-}}$ collisions. In particular, in most calculations [152], a similar mass is obtained or assumed for S=1 (ud,uu,dd) diquarks as contained in $\Sigma_{\mathrm{c}}^{0,+,++}$ and S=0 (ds) diquarks as in $\Xi_{\mathrm{c}}^{0}$. As discussed for the $\Lambda_{\mathrm{c}}^{+}\leftarrow\Sigma_{\mathrm{c}}^{0,+,++}/\Lambda_{\mathrm{c}}^{+}$ ratio, this could suggest that light diquarks are more easily produced in pp than $\mathrm{e^{+}e^{-}}$ collisions via other processes than the Schwinger mechanism (e.g. colour reconnections or coalescence), or that in pp collisions charm-baryon formation does not proceed via the combination of a charm quark with a light diquark. As mentioned in Sect. II.2.4, the coalescence mechanism can be factorized in terms of a formation of diquark recombining with a further quark. In the limit of negligible binding energy this would correspond to the coalescence model already discussed. However, this is different from the assumption of the existence of a thermal yield of diquarks with masses that are not simply the sum of the two quark masses. In this respect it is also argued in Ref. [194] that a further enhancement of $\Xi_{\mathrm{c}}^{+,0}$/$\mathrm{D^{0}}$ ratio could be due to the large attraction of the $us$ and $ds$ diquark channel with respect to the $ud$ one which is suggested by a relativistic quark model with a potential having a large screened mass as extracted from recent LQCD data. Introducing such an additional coalescence component in the Catania model could strongly improve its agreement with experimental data [194], but within the current experimental data uncertainties no conclusion can be drawn. Clarifying the role of light and heavy diquarks will be important for understanding hadronization in pp as well as in heavy-ion collisions, where diquarks could also be effective degrees of freedom of the system close to the phase transition [111].